### v1

# Cubic polynomials and sums of two squares

Abstract.
[…]
Key words and phrases:
Sums of two squares, Cubic Polynomials, Châtelet surface, Diophantine Equation, Units
2020 Mathematics Subject Classification:
11D25, 11D45, 11R09, 11R16, 11R27
1. Introduction
[…]
Theorem 1.1.
If $\mathbf{p}(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ is an irreducible cubic polynomial with $a_{2}^{2}-a_{1}$ even, and $a_{1}a_{2}-a_{0}$ odd, then
$\#B_{\mathbf{p}}(X)\gg X^{1/3-o(1)}.$
Our result builds upon work of Friedlander and Iwaniec regarding values of quadratic polynomials expressible as a value of a quadratic form [2].
Our result is also related to a result of Iwaniec and Munshi [4] regarding densities of rational points on the surface $\phi(y,z)=\mathbf{p}(x)$, where $\phi$ is an irreducible quadratic form with rational coefficients, and $\mathbf{p}(x)$ is an irreducible cubic polynomial in ${\mathbb{Q}}[x]$.
We note that the bound in Theorem 1.1 is not always optimal. If we can find an integer-valued quadratic polynomial $q_{0}$, and polynomials $q_{1},q_{2}\in{\mathbb{Q}}[x]$ such that $\mathbf{p}(q_{0}(x))=q_{1}(x)^{2}+q_{2}(x)^{2}$ then $B_{\mathbf{p}}(X)\gg X^{1/2}$ (this idea was inspired by an online post on MathOverflow by Grechuk [3]). Through trial and error we obtain the identity
$(x^{2}+8)^{3}+17=(x^{3}+10x)^{2}+(2x^{2}+23)^{2},$
so that $B_{x^{3}+17}(X)\gg X^{1/2}$. Also, when $n$ is an integer we trivially have $B_{x^{3}+n^{2}}(X)\gg X^{1/2}$ as for any integer $m$, the quantity $m^{6}+n^{2}$ is a sum of two squares. In general, the author has not been able to find such polynomial substitutions for an arbitrary monic polynomial in ${\mathbb{Z}}[x]$ of degree $3$. We make the following conjecture.
Conjecture 1.2.
Is it true that for any irreducible monic polynomial in ${\mathbb{Z}}[x]$ of degree $3$, say $\mathbf{p}(x)$, there exists an integer-valued quadratic polynomial $q_{0}(x)$, and polynomials $q_{1}(x),q_{2}(x)\in{\mathbb{Q}}[x]$ such that $\mathbf{p}(q_{0}(x))=q_{1}(x)^{2}+q_{2}(x)^{2}$?
Related to this conjecture, a result by Schinzel in [5, Lemma 10] demonstrates that for any such polynomial $\mathbf{p}(x)$, there exists a quadratic polynomial $\mathbf{h}(x)\in{\mathbb{Z}}[x]$ and a cubic polynomial $\mathbf{g}(x)\in{\mathbb{Z}}[x]$ such that $\mathbf{g}(x)$ divides $\mathbf{p}(\mathbf{h}(x))$. Perhaps applying the techniques used in the paper [5] to the polynomial ring ${\mathbb{Z}}[i][x]$ may settle the conjecture.
[…]
We have the following lemma:
Lemma 1.3.
If an integer $z\in{\mathbb{Z}}$ satisfies
[…]
With the ring ${\mathbb{Z}}[\theta]$ we define a notion of maximum coefficient, where $M_{\theta}(c_{0}+c_{1}\theta+c_{2}\theta^{2})=\max\{|c_{0}|,|c_{1}|,|c_{2}|\}$. For $Y,Y_{1},Y_{2}\geq 0$, define
[…]
and
[…]
With Lemma 1.3 we have
(1.1) $\#B_{\mathbf{p}}(X)\geq\#\mathcal{E}_{\theta,q}(Y,X)$
for all $X,Y\geq 0$ and $q\in{\mathbb{N}}$. Let
$\mathcal{R}_{\theta,q}(Y_{1},Y_{2}):=\left\{(\omega_{1},\omega_{2}):\ \begin{aligned} &\omega_{i}\in{\mathbb{Z}}[\theta],\ M_{\theta}(\omega_{i})\leq Y_{1},\\ &\exists n\in{\mathbb{Z}},\ |n|\leq Y_{2},\ \omega_{1}^{2}+\omega_{2}^{2}=q^{2}(n-\theta)\end{aligned}\right\}.$
We relate the sets $\mathcal{R}_{\theta,q}(Y_{1},Y_{2})$ and $\mathcal{E}_{\theta,q}(Y_{1},Y_{2})$.
Lemma 1.4.
If $\theta$ is a real algebraic integer of degree 3, $\varepsilon>0$ and $Y_{1},Y_{2}\geq 2$ we have
$\#\mathcal{E}_{\theta,q}(Y_{1},Y_{2})\gg\#\mathcal{R}_{\theta,q}(Y_{1},Y_{2})\cdot Y_{2}^{-\varepsilon}\cdot\log^{-2}Y_{1},$
where the implied constant depends only on $q,\varepsilon$ and $\theta$.
Lower bounds for $\mathcal{R}_{\theta,q}(\cdot,\cdot)$ alongside Lemma 1.4 and inequality (1.1), can be used to obtain lower-bounds for $B_{\mathbf{p}}(\cdot)$. In the subsequent lemma we pursue lower-bounds for $\mathcal{R}_{\theta,1}(\cdot,\cdot)$ under parity conditions on the coefficients of $\mathbf{p}$.
Lemma 1.5.
If $\mathbf{p}(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ is an irreducible cubic polynomial with $a_{2}^{2}-a_{1}$ even, and $a_{1}a_{2}-a_{0}$ odd, and $\theta$ is a root of $\mathbf{p}$, we have the estimate
$\#\mathcal{R}_{\theta,1}(X,X)\gg X^{1/3-o(1)}.$
Thus Theorem 1.1 follows from Lemma 1.5, Lemma 1.4 and inequality (1.1).
2. Proof of Lemma 1.3
[…]
$z-\theta=(v_{0}+v_{1}\theta+v_{2}\theta^{2})^{2}+(w_{0}+w_{1}\theta+w_{2}\theta^{2})^{2}$
(2.1) $=(v_{0}+v_{1}\theta+v_{2}\theta^{2}+i(w_{0}+w_{1}\theta+w_{2}\theta^{2}))\times$
$\times(v_{0}+v_{1}\theta+v_{2}\theta^{2}-i(w_{0}+w_{1}\theta+w_{2}\theta^{2})),$
for some rational values $v_{d},w_{d}$. Furthermore, whenever $\alpha$ is a root of $\mathbf{p}$, identity (2) holds with $\theta$ replaced with $\alpha$ (due to the isomorphism ${\mathbb{Q}}[i][\theta]\rightarrow{\mathbb{Q}}[i][\alpha]$ fixing ${\mathbb{Q}}[i]$ and sending $\theta$ to $\alpha$). In particular we have that
$P_{1}:=\prod_{\theta:\mathbf{p}(\theta)=0}(v_{0}+v_{1}\theta+v_{2}\theta^{2}+i(w_{0}+w_{1}\theta+w_{2}\theta^{2}))$
is a symmetric polynomial over $\theta_{1},\theta_{2},\theta_{3}$ in ${\mathbb{Q}}[i][\theta_{1},\theta_{2},\theta_{3}]$, where $\theta_{1},\theta_{2},\theta_{3}$ are roots of $\mathbf{p}(x)$ (here we used that $\mathbf{p}(x)$ is irreducible over ${\mathbb{Q}}[i]$). Thus $P_{1}$ is a Gaussian rational, say $u+vi\in{\mathbb{Q}}[i]$. Also note that
$P_{2}$ $:=\prod_{\theta:\ L(\theta)=0}(v_{0}+v_{1}\theta+v_{2}\theta^{2}-i(w_{0}+w_{1}\theta+w_{2}\theta^{2}))$
$=\overline{P_{1}}=u-vi.$
Now with (2) we write
$\mathbf{p}(z)=\prod_{\theta:\ \mathbf{p}(\theta)=0}(z-\theta)=P_{1}P_{2}=u^{2}+v^{2}.$
[…]
3. Proof of Lemma 1.4
For $Y,Y_{1},Y_{2}\geq 0$ and $q\in\mathbb{N}$, define
$\mathcal{A}_{\theta}(\omega,Y):=\left\{(\omega_{1},\omega_{2}):\ \omega_{1}^{2}+\omega_{2}^{2}=\omega,\ M_{\theta}(\omega_{i})\leq Y,\ \omega_{i}\in{\mathbb{Z}}[\theta]\right\},$
and
$\mathcal{C}_{\theta}(Y):=\left\{(c_{0}+id_{0})+(c_{1}+id_{1})\theta+(c_{2}+id_{2})\theta^{2}:\ c_{j},d_{j}\in{\mathbb{Z}},\ |c_{j}|,|d_{j}|\leq Y\right\}.$
Observe that we have
(3.1) $\mathcal{E}_{\theta,q}(Y_{1},Y_{2})=\left\{n:\ n\in{\mathbb{Z}},\ \#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})>0,\ |n|\leq Y_{2}\right\}.$
and let $\mathbf{W}_{\theta}:{\mathbb{Q}}[\theta]\times{\mathbb{Q}}[\theta]\rightarrow{\mathbb{Q}}[\theta,i]$ be the bijective map
$\mathbf{W}_{\theta}(\omega_{1},\omega_{2})=\omega_{1}+i\omega_{2}.$
When $\theta_{1}$ and $\theta_{2}$ are roots of a degree $3$ monic irreducible polynomial, say $p(x)$ in ${\mathbb{Z}}[x]$, we observe that there exists a unique field isomorphism $\kappa_{\theta_{1},\theta_{2}}:{\mathbb{Q}}[\theta_{1},i]\rightarrow{\mathbb{Q}}[\theta_{2},i]$ which fixes ${\mathbb{Q}}[i]$ but sends $\theta_{1}$ to $\theta_{2}$ (as the polynomial $\mathbf{p}(x)$ is irreducible over ${\mathbb{Q}}[i]$).
Lemma 3.1.
Let
$\mathbf{m}(\theta,q,Y_{1},Y_{2})=\max\limits_{n\in\mathbb{Z},\ \lvert n\rvert\leq Y_{2}}\#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})+1.$
We have
$\#\mathcal{E}_{\theta,q}(Y_{1},Y_{2})\geq\frac{\#\mathcal{R}_{\theta,q}(Y_{1},Y_{2})}{\mathbf{m}(\theta,q,Y_{1},Y_{2})}.$
Proof.
Observe that
$\bigcup_{n\in{\mathbb{Z}},\ |n|\leq Y_{2}}\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})=$
$=\left\{(\omega_{1},\omega_{2}):\ \begin{aligned} &\omega_{i}\in{\mathbb{Z}}[\theta],\ M_{\theta}(\omega_{i})\leq Y_{1},\\ &\exists n\in{\mathbb{Z}},\ |n|\leq Y_{2},\ \omega_{1}^{2}+\omega_{2}^{2}=q^{2}(n-\theta)\end{aligned}\right\}$
$=\mathcal{R}_{\theta,q}(Y_{1},Y_{2}).$
Furthermore, for $n_{1}\neq n_{2}$ we have
$\mathcal{A}_{\theta}(q^{2}(n_{1}-\theta),Y_{1})\cap\mathcal{A}_{\theta}(q^{2}(n_{2}-\theta),Y_{1})=\emptyset.$
Hence with (3.1) we have
$\#\mathcal{R}_{\theta,q}(Y_{1},Y_{2})$ $=\sum_{n\in{\mathbb{Z}},\ |n|\leq Y_{2}}\#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})$
$=\sum_{n\in\mathcal{E}_{\theta,q}(Y_{1},Y_{2})}\#\mathcal{A}_{\theta}(q^{2}(n-\theta),Y_{1})$
$\leq\mathbf{m}(\theta,q,Y_{1},Y_{2})\cdot\#\mathcal{E}_{\theta,q}(Y_{1},Y_{2}).$
$\sqcap\sqcup$
Lemma 3.2.
If $Y\geq 1$ and $w\in\mathcal{C}_{\theta}(Y)$ is a non-zero element of ${\mathbb{Q}}[\theta,i]$ then
$Y^{-2}\ll|w|\ll Y,$
where $|\cdot|$ denotes the complex modulus operator.
Proof.
The upper-bound follows from the triangle inequality. Let $\mathbf{p}(x)$ be the degree $3$ monic irreducible polynomial for $\theta$ with roots $\theta_{1},\theta_{2}$ and $\theta_{3}$ respectively, where without loss of generality we may assume $\theta=\theta_{1}$. Suppose that
$w=(c_{0}+id_{0})+(c_{1}+id_{1})\theta+(c_{2}+id_{2})\theta^{2}$
for integers $c_{j},d_{j}$, then
$\prod_{j=1}^{3}\left((c_{0}+id_{0})+(c_{1}+id_{1})\theta_{j}+(c_{2}+id_{2})\theta_{j}^{2}\right)$
is a symmetric polynomial in ${\mathbb{Z}}[i][\theta_{1},\theta_{2},\theta_{3}]$ and thus is a guassian integer, necessarily non-zero as $w\neq 0$. Thus we have
$|w|\left|\prod_{j=1}^{2}\left((c_{0}+id_{0})+(c_{1}+id_{1})\theta_{j}+(c_{2}+id_{2})\theta_{j}^{2}\right)\right|\geq 1,$
or $|w|\cdot Y^{2}\gg 1$.   $\sqcap\sqcup$
Lemma 3.3.
If $\beta$ is an algebraic integer of degree $3$, then ${\mathrm{Im}}(\beta)\not\in{\mathbb{Z}}\setminus\{0\}$.
Proof.
Suppose $\beta=\mu+ih$ is an algebraic integer of degree $3$ for some $h\neq 0,h\in{\mathbb{Z}}$ and $\mu\in{\mathbb{R}}$. Then it has a minimal polynomial $x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$ with roots $\mu+ih$, $\mu-ih,$ and $\theta$, where $\theta$ is real. By comparing with values of coefficients we have $\theta+2\mu\in{\mathbb{Z}}$ and $2\theta\mu+\mu^{2}+h^{2}\in{\mathbb{Z}}$, therefore $\theta$ is a root of a quadratic polynomial in ${\mathbb{Q}}[x]$, but this is impossible as $\theta$ is of degree $3$ (otherwise $\beta$ will be of degree $\leq 2$).   $\sqcap\sqcup$
Lemma 3.4.
If $\theta$ is an algebraic integer of degree $3$ with minimal polynomial $P(x)$ then $\theta+i$ is an algebraic integer of degree $6$. Furthermore the minimal polynomial of $\theta+i$ over $\mathbb{Z}$} is $Q(x)$, where
$Q(x)=\prod_{d=1}^{3}(x-(\theta_{d}+i))(x-(\theta_{d}-i)),$
here $\theta_{1},\theta_{2}$ and $\theta_{3}$ are the distinct roots of $P(x)$.
Proof.
Let $P(x)=x^{3}+a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{Z}}[x]$. We can consider the polynomials $P_{1}(x)$ and $P_{2}(x)$ in ${\mathbb{Z}}[i][x]$ which are defined as the binomial expansions of $P(x+i)$ and P(xi)P(x-i) respectively. It is clear that $P(x)$ is irreducible over ${\mathbb{Q}}[i]$ (otherwise ${\mathbb{Q}}[i]:{\mathbb{Q}}]\geq 3$ which is impossible), thus $P_{1}(x)$ and $P_{2}(x)$ are irreducible over ${\mathbb{Q}}[i]$. Both $P_{1}(x)$ and $P_{2}(x)$ are not polynomials in ${\mathbb{Z}}[x]$ as the $x^{2}$ coefficients of the polynomials are $3i+a_{2}$ and $(-3i+a_{2})$ respectively. Observe that the polynomial $Q(x)=P_{1}(x)P_{2}(x)$ has integer valued coefficients, thus the minimal polynomial of $\theta+i$ over the integers divides $Q(x)$.
If $Q(x)$ factors as $H_{1}(x)H_{2}(x)$, where $H_{j}(x)\in\mathbb{Q}[x]$ and $\deg(H_{j}(x))\geq 1$, then it factors similarly as $H_{1}(x)\cdot H_{2}(x)$ in ${\mathbb{Q}}[i][x]$. Since ${\mathbb{Q}}[i][x]$ is a unique factorization domain with units $c$, where $c\in{\mathbb{Q}}[i]\setminus\{0\}$, we have $H_{1}(x)=cP_{\sigma(1)}(x)$ and $H_{2}(x)=(1/c)P_{\sigma(2)}(x)$, where $\sigma:\{1,2\}\rightarrow\{1,2\}$ is a permutation map and $c\in{\mathbb{Q}}[i]$, $c\neq 0$. If $c$ is real then $H_{j}(x)\not\in{\mathbb{Z}}[x]$, which is impossible. If $c$ is complex then the $x^{3}$ coefficients of $H_{1}(x),H_{2}(x)$ are complex which is also impossible. Hence the polynomial $Q(x)$ is irreducible over $\mathbb{Z}$}.   $\sqcap\sqcup$
Lemma 3.5.
If $\theta$ is an algebraic integer of degree $3$ then the group of units of $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ has rank of two.
Proof.
If $K={\mathbb{Q}}[\theta,i]$, then by Lemma 3.4 we have $K={\mathbb{Q}}[\theta+i]$, with the minimal polynomial of $\theta+i$ being $Q(x)$. By Lemma 3.3, all the roots of $Q$ are complex-valued. Thus by Dirichlet’s unit theorem the rank of the group of units of $\mathcal{O}_{K}$ is (6/2)1=2(6/2)-1=2.   $\sqcap\sqcup$
Lemma 3.6.
If $\theta$ is a real algebraic integer of degree $3$, there exists a finite set $W$ of units in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$, such that for every unit $u\in\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ there exists integers $t_{1}$ and $t_{2}$ so that
$u=w\cdot l_{1}^{t_{1}}l_{2}^{t_{2}},$
where $l_{1},l_{2}$ are units in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ with the complex modulus of $l_{1}$ not equal to 11 ($|l_{1}|\neq 1$), $|l_{2}|=1$, and $w$ is some unit belonging to $W$.
Proof.
By lemma 3.5, there exists two multiplicatively independent units $u_{1}$ and $u_{2}$ in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ so that every unit $u$ in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ is of the form
$u=T\cdot u_{1}^{t_{1}}u_{2}^{t_{2}},$
for some $T\in\text{Tor}(\mathcal{O}_{{\mathbb{Q}}[\theta,i]})$ and integers $t_{1}$ and $t_{2}$. By Dirichlet’s unit theorem there exists a unit $q_{1}$ of infinite order in $\mathcal{O}_{{\mathbb{Q}}[\theta]}$, which is also a real unit in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$. We necessarily have $|q_{1}|\neq 1$.
Since ${\mathbb{Q}}[\theta,i]$ is closed under complex conjugation, by [1, Theorem 1] there exists a unit $q_{2}$ of $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ of infinite order with $|q_{2}|=1$ (this unit is not a torsion element). We observe that there exists integers $s_{1},s_{2},y_{1},y_{2}$ and two torsion elements $g,h\in\text{Tor}(\mathcal{O}_{{\mathbb{Q}}[\theta,i]})$ so that
$q_{1}=g\cdot u_{1}^{s_{1}}u_{2}^{s_{2}}$
and
$q_{2}=h\cdot u_{1}^{y_{1}}u_{2}^{y_{2}}.$
Take $l_{1}=q_{1}g^{-1}$ and $l_{2}=q_{2}h^{-1}$. Observe that $l_{1}$ and $l_{2}$ are multiplicatively independent, thus the matrix
$\begin{bmatrix}s_{1}&y_{1}\\ s_{2}&y_{2}\end{bmatrix}$
has non-zero determinant. Let $D=|s_{1}y_{2}-y_{1}s_{2}|$, so that $D\geq 1$, and put
$W=\{T\cdot u_{1}^{v_{1}}u_{2}^{v_{2}}:\ T\in\text{Tor}(\mathcal{O}_{{\mathbb{Q}}[\theta,i]}),\ v_{i}\in\{1,\ldots,D\}\}.$
The lemma follows with the choices of $l_{1},l_{2}$ and $W$. Here it is useful to recall that $\text{Tor}(\mathcal{O}_{{\mathbb{Q}}[\theta,i]})<\infty$, and the torsion points of $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ are precisely the roots of cyclotomic polynomials (which have complex modulus one) within ${\mathbb{Q}}[\theta,i]$.   $\sqcap\sqcup$
Lemma 3.7.
If $\theta$ is a real algebraic integer of degree $3$, $l_{2}\in\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ is a unit of infinite order with $|l_{2}|=1$, $b\in{\mathbb{Q}}[\theta,i]$ is non-zero, and $Y\geq 2$ then
$\#\left\{n:n\in{\mathbb{Z}},\ b\cdot l_{2}^{n}\in\mathcal{C}_{\theta}(Y)\right\}\ll\log Y,$
where the implied constant is independent of $b$, but dependent on $l_{2}$.
Proof.
Let $p(x)$ be the minimal monic polynomial of $\theta$ in ${\mathbb{Z}}[x]$, which is necessarily irreducible over ${\mathbb{Q}}[i]$; also let $\theta_{1},\theta_{2},\theta_{3}$ be the roots of $p(x)$ with $\theta=\theta_{1}$. Observe that $l_{2}$ is the root of the polynomial $q(x)$ where
$q(x)=\prod_{j=1}^{3}\left(x-\kappa_{\theta_{1},\theta_{j}}(l_{2})\right)\left(x-\overline{\kappa_{\theta_{1},\theta_{j}}(l_{2})}\right).$
Observe that if $x$ is real then $q(x)$ is real, thus $q(x)\in{\mathbb{R}}[x]$. Since $p(x)$ is irreducible over ${\mathbb{Q}}[i]$, the polynomials $\prod_{j=1}^{3}\left(x-\kappa_{\theta_{1},\theta_{j}}(l_{2})\right)$ and $\prod_{j=1}^{3}\left(x-\overline{\kappa_{\theta_{1},\theta_{j}}(l_{2})}\right)$ belong to ${\mathbb{Q}}[i]$ (as both expressions can be expanded as a symmetric polynomial in $\theta_{1},\theta_{2},\theta_{3}$). Hence we deduce that $q(x)\in{\mathbb{R}}[x]\cap\left({\mathbb{Q}}[i][x]\right)$, implying that $q(x)\in{\mathbb{Q}}[x]$. By Kronecker’s theorem there exists a choice of $j\in\{2,3\}$ so that $\left|\kappa_{\theta_{1},\theta_{j}}(l_{2})\right|\neq 1$. Now note that we wish to bound from above the cardinality of the set $\mathcal{X}_{\theta}(b,l_{2},Y)$, where
$\mathcal{X}_{\theta}(b,l_{2},Y)=\{n:n\in{\mathbb{Z}},\ b\cdot l_{2}^{n}\in\mathcal{C}_{\theta}(Y)\}.$
Observe that
$\mathcal{X}_{\theta}(b,l_{2};Y)=\mathcal{X}_{\theta_{j}}(\kappa_{\theta_{1},\theta_{j}}(b),\kappa_{\theta_{1},\theta_{2}}(l_{2});Y).$
Let $\rho=|\kappa_{\theta_{1},\theta_{2}}(l_{2})|$ and $c=|\kappa_{\theta_{1},\theta_{2}}(b)|$. By lemma 3.2 there exists constants $A,B>0$ depending on $\theta$ so that
$\mathcal{X}_{\theta_{j}}(\kappa_{\theta_{1},\theta_{j}}(b),\kappa_{\theta_{1},\theta_{2}}(l_{2});Y)\subseteq\{n:\ n\in{\mathbb{Z}},\ A\cdot Y^{-2}\leq c\cdot\rho^{n}\leq B\cdot Y\}.$
The lemma now follows (here we use $\rho\neq 1$).   $\sqcap\sqcup$
For a non-zero element $w\in{\mathbb{Z}}[i,\theta]$ let $\mathcal{U}(w,Y)$ denote elements $s$ of $\mathcal{C}_{\theta}(Y)$ where $s=wu$ for some unit $u$ in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$.
Lemma 3.8.
If $\theta$ is a real algebraic integer of degree 3, $w\neq 0$ is an element of ${\mathbb{Z}}[\theta,i]$ and $Y\geq 2$, we have
$\#\mathcal{U}(w,Y)\ll\log^{2}Y,$
where the implied constant is independent of $w$.
Proof.
From lemma 3.6, there exists a finite set $W$ and units $l_{1}$ and $l_{2}$ in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ which satisfy $|l_{1}|\neq 1$ and $|l_{2}|=1$ so that for every unit $u$ in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ there exists $g\in W$ and integers $t_{1}$ and $t_{2}$ with
$u=g\cdot l_{1}^{t_{1}}l_{2}^{t_{2}}.$
We may now partition the set $\mathcal{U}(w,Y)$ into sets $\mathcal{U}(g,t_{1};w,Y)$ where $g\in W$ and $t_{1}$ is an integer, more specifically, this defined subset satisfies
$\mathcal{U}(g,t_{1};w,Y)=\{t_{2}:\ t_{2}\in{\mathbb{Z}},\ wgl_{1}^{t_{1}}l_{2}^{t_{2}}\in\mathcal{C}_{\theta}(Y)\}.$
By lemma 3.2 there exists positive constants $A$ and $B$ (depending only on $\theta$ and the set $W$) so that when $|w||g||l_{1}|^{t_{1}}>B\cdot Y$ or $|w||g||l_{1}|^{t_{1}}<A\cdot Y^{-2}$, the set $\mathcal{U}(g,t_{1};w,Y)$ is empty. Define the quantities
$H(g)=\left\lceil\log_{|l_{1}|}\left(\frac{BY}{|w||g|}\right)\right\rceil$
$h(g)=\left\lfloor\log_{|l_{1}|}\left(\frac{AY^{-2}}{|w||g|}\right)\right\rfloor$
(here the logarithm is taken in base $|l_{1}|$). We observe
$\mathcal{U}(w,Y)=\bigcup_{g\in W}\bigcup_{t_{1}=h(g)}^{H(g)}\mathcal{U}(g,t_{1};w,Y).$
Hence by lemma 3.7 we have
$\#\mathcal{U}(w,Y)$ $\ll\sum_{g\in W}\left|H(g)-h(g)+1\right|\cdot\log Y$
$\ll\log^{2}Y.$
$\sqcap\sqcup$
Lemma 3.9.
If $\theta$ is a real algebraic integer of degree $3$, and $\omega\in{\mathbb{Z}}[\theta]$. Then for every $\varepsilon>0$ and $Y\geq 2$ we have
$\#\mathcal{A}_{\theta}(\omega,Y)\ll M(\omega)^{\varepsilon}\cdot\log^{2}Y,$
where the implied constant depends only on $\theta$ and $\varepsilon$ (effective only with sufficient knowledge of units in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$).
Proof.
The bound is trivial for $\omega=0$; for the rest of the proof assume that $\omega\neq 0$. Define the set $d(\omega)$ to contain ideals in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ that divide the ideal $(\omega)$. Since $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$ is a Dedekind domain, we have
$\#d(\omega)\ll N_{{\mathbb{Q}}[\theta,i]/{\mathbb{Q}}}(\omega)^{\varepsilon}\ll M(\omega)^{6\varepsilon}.$
Also let $p(\omega)$ denote the principal ideals of $d(\omega)$ so that $p(\omega)\subset d(\omega)$. We can now define the map $I_{\omega,Y}:\mathcal{A}_{\theta}(\omega,Y)\rightarrow p(\omega)$ where
$I_{\omega,Y}(\omega_{1},\omega_{2})=(\omega_{1}+i\omega_{2}).$
The map $I_{\omega,Y}$ induces an equivalence relation on $\mathcal{A}_{\theta}(\omega,Y)$ where $\mathbf{a}\sim\mathbf{b}$ if $I_{\omega,Y}(\mathbf{a})=I_{\omega,Y}(\mathbf{b})$. Accordingly we may partition $\mathcal{A}_{\theta}(\omega,Y)$ into $t$ equivalence classes, where $t\ll M(\omega)^{6\varepsilon}$. Pick elements $(u_{1},v_{1}),\ldots(u_{t},v_{t})$ of $\mathcal{A}_{\theta}(\omega,Y)$ which belong to distinct equivalence classes. Define
$\mathcal{A}_{\theta}(u_{j},v_{j};\omega,Y)=\{\mathbf{s}:\ \mathbf{s}\in\mathcal{A}_{\theta}(\omega,Y),\ \mathbf{s}\sim(u_{j},v_{j})\}$
so that
(3.2) $\bigcup_{j=1}^{t}\mathcal{A}_{\theta}(u_{j},v_{j};\omega,Y)=\mathcal{A}_{\theta}(\omega,Y).$
We now have to estimate the cardinality of $\mathcal{A}_{\theta}(u_{j},v_{j};\omega,Y)$. Observe that if $(\alpha,\beta)\in\mathcal{A}_{\theta}(u_{j},v_{j};\omega,Y)$ then $(\alpha+i\beta)=(u_{j}+iv_{j})$ (as ideals in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$) and $\alpha+i\beta\in\mathcal{C}_{\theta}(Y)$. The ideal equality requirement happens if and only if $\alpha+i\beta=(u_{j}+iv_{j})u$ for some unit $u$ in $\mathcal{O}_{{\mathbb{Q}}[\theta,i]}$. Hence we have
$\mathbf{W}_{\theta}(\mathcal{A}_{\theta}(u_{j},v_{j};\omega,Y))=\mathcal{U}_{\theta}(u_{j}+iv_{j},Y).$
By lemma 3.8 we have $\mathcal{U}_{\theta}(u_{j}+iv_{j},Y)\ll\log^{2}Y$. Thus
$\#\mathcal{A}_{\theta}(u_{j},v_{j};\omega,Y)=\#\mathbf{W}_{\theta}(\mathcal{A}_{\theta}(u_{j},v_{j};\omega,Y))\ll\log^{2}Y.$
By the set relation (3.2) we have
$\#\mathcal{A}_{\theta}(\omega,Y)$ $\ll t\cdot\log^{2}Y$
$\ll M(\omega)^{6\varepsilon}\cdot\log^{2}Y.$
$\sqcap\sqcup$
We now remark that Lemma 1.4 follows from Lemma 3.1 and Lemma 3.9.
4. Proof of Lemma 1.5
Throughout this section assume that $u_{0},u_{1},u_{2},v_{0},v_{1},v_{2}$ are integer- valued variables. We also abbreviate $(u_{0},u_{1},u_{2},v_{0},v_{1},v_{2})$ as $(\mathbf{u},\mathbf{v})$. Define the following functions:
$g_{0}(\mathbf{u},\mathbf{v})=$ $(u_{0}^{2}-2a_{0}u_{1}u_{2}+u_{2}^{2}a_{0}a_{2})+$
$+(v_{0}^{2}-2a_{0}v_{1}v_{2}+v_{2}^{2}a_{0}a_{2}),$
$g_{1}(\mathbf{u},\mathbf{v})=$ $\left(2u_{0}u_{1}-2a_{1}u_{1}u_{2}+u_{2}^{2}(a_{1}a_{2}-a_{0})\right)+$
$+\left(2v_{0}v_{1}-2a_{1}v_{1}v_{2}+v_{2}^{2}(a_{1}a_{2}-a_{0})\right),$
$g_{2}(\mathbf{u},\mathbf{v})=$ $(2u_{0}u_{2}+u_{1}^{2}-2a_{2}u_{1}u_{2}+u_{2}^{2}(a_{2}^{2}-a_{1}))+$
$+(2v_{0}v_{2}+v_{1}^{2}-2a_{2}v_{1}v_{2}+v_{2}^{2}(a_{2}^{2}-a_{1})).$
We have the following lemma:
Lemma 4.1.
We have
$(u_{0}+u_{1}\theta+u_{2}\theta^{2})^{2}+(v_{0}+v_{1}\theta+v_{2}\theta^{2})^{2}=$
$g_{0}(\mathbf{u},\mathbf{v})+g_{1}(\mathbf{u},\mathbf{v})\theta+g_{2}(\mathbf{u},\mathbf{v})\theta^{2}.$
Proof.
We have
$\theta^{3}=-a_{2}\theta^{2}-a_{1}\theta-a_{0},$
$\theta^{4}$ $=-a_{2}\theta^{3}-a_{1}\theta^{2}-a_{0}\theta$
$=-a_{2}(-a_{2}\theta^{2}-a_{1}\theta-a_{0})-a_{1}\theta^{2}-a_{0}\theta$
$=(a_{2}^{2}-a_{1})\theta^{2}+(a_{1}a_{2}-a_{0})\theta+a_{0}a_{2}.$
We expand
$(u_{0}+u_{1}\theta+u_{2}\theta^{2})^{2}=u_{0}^{2}+2u_{0}u_{1}\theta+2u_{0}u_{2}\theta^{2}+u_{1}^{2}\theta^{2}+2u_{1}u_{2}\theta^{3}+u_{2}^{2}\theta^{4}$
$=(u_{0}^{2}-2a_{0}u_{1}u_{2})+(2u_{0}u_{1}-2a_{1}u_{1}u_{2})\theta+(2u_{0}u_{2}+u_{1}^{2}-2a_{2}u_{1}u_{2})\theta^{2}+u_{2}^{2}\theta^{4}$
$=(u_{0}^{2}-2a_{0}u_{1}u_{2})+(2u_{0}u_{1}-2a_{1}u_{1}u_{2})\theta+(2u_{0}u_{2}+u_{1}^{2}-2a_{2}u_{1}u_{2})\theta^{2}+$
$+u_{2}^{2}a_{0}a_{2}+u_{2}^{2}(a_{1}a_{2}-a_{0})\theta+u_{2}^{2}(a_{2}^{2}-a_{1})\theta^{2}$
$=(u_{0}^{2}-2a_{0}u_{1}u_{2}+u_{2}^{2}a_{0}a_{2})+(2u_{0}u_{1}-2a_{1}u_{1}u_{2}+u_{2}^{2}(a_{1}a_{2}-a_{0}))\theta+$
$+(2u_{0}u_{2}+u_{1}^{2}-2a_{2}u_{1}u_{2}+u_{2}^{2}(a_{2}^{2}-a_{1}))\theta^{2}.$
We similarly expand $(v_{0}+v_{1}\theta+v_{2}\theta^{2})^{2}$ and verify the lemma.   $\sqcap\sqcup$
Define
$I(u_{1},u_{2},v_{1},v_{2})=u_{1}v_{2}-v_{1}u_{2},$
$h_{1}(u_{1},u_{2},$ $v_{1},v_{2})=$
$-1+2a_{1}u_{1}u_{2}-u_{2}^{2}(a_{1}a_{2}-a_{0})+2a_{1}v_{1}v_{2}-v_{2}^{2}(a_{1}a_{2}-a_{0}),$
$h_{2}(u_{1},u_{2},$ $v_{1},v_{2})=$
$-u_{1}^{2}+2a_{2}u_{1}u_{2}-u_{2}^{2}(a_{2}^{2}-a_{1})-v_{1}^{2}+2a_{2}v_{1}v_{2}-v_{2}^{2}(a_{2}^{2}-a_{1}),$
$U_{0}(u_{1},u_{2},v_{1},v_{2})=v_{2}h_{1}(u_{1},u_{2},v_{1},v_{2})/2-v_{1}h_{2}(u_{1},u_{2},v_{1},v_{2})/2$
and
$V_{0}(u_{1},u_{2},v_{1},v_{2})=-u_{2}h_{1}(u_{1},u_{2},v_{1},v_{2})/2+u_{1}h_{2}(u_{1},u_{2},v_{1},v_{2})/2.$
Let $\mathbf{S}\subset{\mathbb{Z}}[\theta]\times{\mathbb{Z}}[\theta]$ contain tuples $((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2}))$ such that
  • 1) $v_{2}$ is even, $\gcd(u_{2},v_{2})=1$
  • 2) $u_{1}$ and $v_{1}$ are odd, with $I(u_{1},u_{2},v_{1},v_{2})=1$
  • 3) $u_{0}=U_{0}(u_{1},u_{2},v_{1},v_{2})$ and $v_{0}=V_{0}(u_{1},u_{2},v_{1},v_{2})$.
Lemma 4.2.
We have $\mathbf{S}\subset\mathcal{R}_{\theta,1}(\infty,\infty)$.
Proof.
By Lemma 4.1 we have
$\left((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2})\right)\in\mathcal{R}_{\theta,1}(\infty,\infty)$
if and only if $g_{1}(\mathbf{u},\mathbf{v})=-1$ and $g_{2}(\mathbf{u},\mathbf{v})=0$. Solving the two equations gives us
$2u_{1}u_{0}+2v_{1}v_{0}=-1+2a_{1}u_{1}u_{2}-u_{2}^{2}(a_{1}a_{2}-a_{0})+2a_{1}v_{1}v_{2}-v_{2}^{2}(a_{1}a_{2}-a_{0}),$
and
$2u_{2}u_{0}+2v_{2}v_{0}=-u_{1}^{2}+2a_{2}u_{1}u_{2}-u_{2}^{2}(a_{2}^{2}-a_{1})-v_{1}^{2}+2a_{2}v_{1}v_{2}-v_{2}^{2}(a_{2}^{2}-a_{1}),$
where the right hand side of the above equations are $h_{1}(u_{1},u_{2},v_{1},v_{2})$ and $h_{2}(u_{1},u_{2},v_{1},v_{2})$ respectively. This leads us to the matrix equation
$\begin{bmatrix}2u_{1}&2v_{1}\\ 2u_{2}&2v_{2}\end{bmatrix}\times\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\begin{bmatrix}h_{1}(u_{1},u_{2},v_{1},v_{2})\\ h_{2}(u_{1},u_{2},v_{1},v_{2})\end{bmatrix}$.
If $I(u_{1},u_{2},v_{1},v_{2})\neq 0$, solving the matrix equation gives us
$\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}$ $=\frac{1}{u_{1}v_{2}-v_{1}u_{2}}\begin{bmatrix}v_{2}&-v_{1}\\ -u_{2}&u_{1}\end{bmatrix}\times\begin{bmatrix}h_{1}(u_{1},u_{2},v_{1},v_{2})/2\\ h_{2}(u_{1},u_{2},v_{1},v_{2})/2\end{bmatrix}$
$=\frac{1}{u_{1}v_{2}-v_{1}u_{2}}\begin{bmatrix}v_{2}h_{1}(u_{1},u_{2},v_{1},v_{2})/2-v_{1}h_{2}(u_{1},u_{2},v_{1},v_{2})/2\\ -u_{2}h_{1}(u_{1},u_{2},v_{1},v_{2})/2+u_{1}h_{2}(u_{1},u_{2},v_{1},v_{2})/2\end{bmatrix}$
$=\frac{1}{I(u_{1},u_{2},v_{1},v_{2})}\begin{bmatrix}U_{0}(u_{1},u_{2},v_{1},v_{2})\\ V_{0}(u_{1},u_{2},v_{1},v_{2}).\end{bmatrix}$
We thus verify that $\mathbf{S}\subset\mathcal{R}_{\theta,1}(\infty,\infty)$.   $\sqcap\sqcup$
We now more closely examine the set $\mathbf{S}$. From now on we will denote $\alpha$ and $\beta$ to be coprime integers such that $\alpha$ is even. Note that there exists odd integers $\mathbf{v}(\alpha,\beta)$ and $\mathbf{u}(\alpha,\beta)$ satisfying $\mathbf{u}(\alpha,\beta)\alpha-\mathbf{v}(\alpha,\beta)\beta=1$, and in particular
$(\mathbf{u}(\alpha,\beta)+2\beta t)\alpha-(\mathbf{v}(\alpha,\beta)+2\alpha t)\beta=1$
for all integers $t$. Observe that $\mathbf{S}$ satisfies a completion property. If integer values $v_{2}=\alpha$, $u_{2}=\beta$ and $v_{1}$ are given with $v_{1}\equiv\mathbf{v}(\alpha,\beta)\pmod{2\alpha}$, then we can find a unique set of integer values $u_{0},v_{0},v_{1}$ so that
$((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2}))\in\mathbf{S}.$
For $c>0$ we define a subset of $\mathbf{S}$ where
$\mathbf{S}_{c}(\alpha,\beta,X)=$
$\left\{((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2})):\ \begin{aligned} &v_{2}=\alpha,\ u_{2}=\beta,\ |v_{1}|\leq c\cdot X^{1/6},\\ &v_{1}\equiv\mathbf{v}(\alpha,\beta)\pmod{2\alpha},\ (\mathbf{u},\mathbf{v})\in\mathbf{S}\end{aligned}\right\}.$
Furthermore define
$\mathbf{V}_{c}(\alpha,\beta,X)=\{v_{1}:\ v_{1}\equiv\mathbf{v}(\alpha,\beta)\pmod{2\alpha},\ |v_{1}|\leq c\cdot X^{1/6}\},$
where by the completion property for $\mathbf{S}$ we remark that the sets $\mathbf{V}_{c}(\alpha,\beta,X)$ and $\mathbf{S}_{c}(\alpha,\beta,X)$ are in bijection and in particular we have
(4.1) $\#\mathbf{V}_{c}(\alpha,\beta,X)=\#\mathbf{S}_{c}(\alpha,\beta,X).$
Lemma 4.3.
If $c>0$, there exists a positive constant $F(c)$ with $\lim_{c\rightarrow 0^{+}}F(c)=0$ so that if $X$ is sufficiently large ($X\geq N_{c}$) and
$1\leq\beta\leq\alpha\leq c\cdot X^{1/6},$
we have
$\mathbf{S}_{c}(\alpha,\beta,X)\subset\mathcal{R}_{\theta,1}(F(c)\cdot X^{1/2},\infty).$
Proof.
If
$((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2}))\in\mathbf{S}_{c}(\alpha,\beta,X),$
by Lemma 4.2 we know that
(4.2) $((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2}))\in\mathcal{R}_{\theta,1}(\infty,\infty).$
It remains to show that $|u_{i}|,|v_{i}|\leq F(c)\cdot X^{1/2}$ for some function $F$ which satisfies $\lim_{c\rightarrow 0^{+}}F(c)=0$. Note that $I(u_{1},\beta,v_{1},\alpha)=1$ or
$u_{1}\alpha-v_{1}\beta=1,$
implying that
$|u_{1}|=\left|\frac{1+v_{1}\beta}{\alpha}\right|\leq 2|v_{1}|.$
Using
(4.3) $|v_{1}|\leq c\cdot X^{1/6},$
we obtain
(4.4) $|u_{1}|\leq 2c\cdot X^{1/6}.$
On the set $\mathbf{S}_{c}(\alpha,\beta,X)$ we thus have
$h_{1}(u_{1},u_{2},v_{1},v_{2})=h_{1}(u_{1},\beta,v_{1},\alpha)$
$=-1+2a_{1}u_{1}\beta-\beta^{2}(a_{1}a_{2}-a_{0})+2a_{1}v_{1}\alpha-\alpha^{2}(a_{1}a_{2}-a_{0}),$
giving us the estimate
$\left|h_{1}(u_{1},u_{2},v_{1},v_{2})\right|\leq 1+2|a_{1}|c^{2}X^{1/3}+2c^{2}|(a_{1}a_{2}-a_{0})|X^{1/3}+2c^{2}|a_{1}|X^{1/3},$
so that if $X$ is sufficiently large we have
$\left|h_{1}(u_{1},u_{2},v_{1},v_{2})\right|$ $\leq\left(4|a_{1}|c^{2}+4c^{2}|(a_{1}a_{2}-a_{0})|+4c^{2}|a_{1}|\right)X^{1/3}$
(4.5) $=H_{1}(c)\cdot X^{1/3},$
where $H_{1}(c)>0$ is some constant that approaches 0 as $c$ approaches 0. Similarly we obtain
$h_{2}(u_{1},u_{2},v_{1},v_{2})=h_{2}(u_{1},\beta,v_{1},\alpha)$
$=-u_{1}^{2}+2a_{2}u_{1}\beta-\beta^{2}(a_{2}^{2}-a_{1})-v_{1}^{2}+2a_{2}\alpha v_{2}-\alpha^{2}(a_{2}^{2}-a_{1}),$
so that with (4.3) and (4.4) we have
(4.6) $|h_{2}(u_{1},u_{2},v_{1},v_{2})|\leq H_{2}(c)\cdot X^{1/3},$
where $H_{2}(c)>0$ is some constant that approaches 0 as $c$ approaches 0. Note that
$u_{0}$ $=U_{0}(u_{1},u_{2},v_{1},v_{2})$
$=v_{2}h_{1}(u_{1},u_{2},v_{1},v_{2})/2-v_{1}h_{2}(u_{1},u_{2},v_{1},v_{2})/2.$
Hence
$|u_{0}|\leq|v_{2}h_{1}(u_{1},u_{2},v_{1},v_{2})|+|v_{1}h_{2}(u_{1},u_{2},v_{1},v_{2})|,$
Using inequalities (4.3), (4.4), (4), (4.6) and substitutions $v_{2}=\alpha$ and $u_{2}=\beta$ we obtain
(4.7) $|u_{0}|\leq\Theta(c)\cdot X^{1/2},$
where $\Theta(c)>0$ is some constant that approaches 0 as $c$ approaches 0. Similarly we obtain
$v_{0}$ $=V_{0}(u_{1},u_{2},v_{1},v_{2})$
$=-u_{2}h_{1}(u_{1},u_{2},v_{1},v_{2})/2+u_{1}h_{2}(u_{1},u_{2},v_{1},v_{2})/2.$
Using inequalities (4.3), (4.4), (4), (4.6) and substitutions $v_{2}=\alpha$ and $u_{2}=\beta$ we obtain
(4.8) $|v_{0}|\leq\rho(c)\cdot X^{1/2},$
where $\rho(c)>0$ is some constant that approaches 0 as $c$ approaches 0. With the choice $F(c)=\max\{2c,\rho(c),\Theta(c)\}$, inequalities (4.3), (4.4), (4.7), (4.8), and set relation (4.2), and the discussion stated immediately after (4.2) the proof follows.   $\sqcap\sqcup$
Lemma 4.4.
If $c>0$ and $2\leq\alpha\leq(c/4)\cdot X^{1/6}$ then
$\#\mathbf{S}_{c}(\alpha,\beta,X)\geq\frac{c\cdot X^{1/6}}{4\alpha}.$
Proof.
By equation (4.1) we have $\mathbf{S}_{c}(\alpha,\beta,X)=\#\mathbf{V}_{c}(\alpha,\beta,X)$, we recall that
$\mathbf{V}_{c}(\alpha,\beta,X)=\{v_{1}:\ v_{1}\equiv\mathbf{v}(\alpha,\beta)\pmod{2\alpha},\ |v_{1}|\leq c\cdot X^{1/6}\}.$
Hence
$\#\mathbf{V}_{c}(\alpha,\beta,X)\geq\left\lfloor\frac{c\cdot X^{1/6}}{2\alpha}\right\rfloor\geq\frac{c\cdot X^{1/6}}{4\alpha}.$
$\sqcap\sqcup$
Lemma 4.5.
There exists a constant $d>0$ so that for sufficiently large $X$ we have
$\mathcal{R}_{\theta,1}(dX^{1/2},\infty)=\mathcal{R}_{\theta,1}(dX^{1/2},X).$
Proof.
Let $d>0$ be sufficiently small so that $2(1+|\theta|+|\theta|^{2})^{2}d^{2}<1$. Observe that if
$((u_{0}+u_{1}\theta+u_{2}\theta^{2}),(v_{0}+v_{1}\theta+v_{2}\theta^{2}))\in\mathcal{R}_{\theta,1}(dX^{1/2},\infty),$
then
$(u_{0}+u_{1}\theta+u_{2}\theta^{2})^{2}+(v_{0}+v_{1}\theta+v_{2}\theta^{2})^{2}=n-\theta$
for some integer $n$. Using triangle inequality we find that
$|n|\leq 2d^{2}(1+|\theta|+|\theta|^{2})^{2}X+|\theta|,$
which is less than $X$ when $X$ is sufficiently large. Hence for large values of $X$ we have
$\mathcal{R}_{\theta,1}(dX^{1/2},\infty)\subseteq\mathcal{R}_{\theta,1}(dX^{1/2},X),$
and the reverse inclusion follows trivially.   $\sqcap\sqcup$
As in Lemma 4.5, choose $d>0$ and a positive real number $N_{d}$ so that for $X\geq N_{d}$ we have
$\mathcal{R}_{\theta,1}(dX^{1/2},\infty)=\mathcal{R}_{\theta,1}(dX^{1/2},X).$
As in Lemma 4.3, choose $c>0$ so that $c<d$ and $F(c)<d$. Observe that if $(\alpha,\beta)\neq(\alpha^{\prime},\beta^{\prime})$ then the sets $\mathbf{S}_{c}(\alpha^{\prime},\beta^{\prime},X)$ and $\mathbf{S}_{c}(\alpha,\beta,X)$ are disjoint. Hence with $\varepsilon>0$, when $X$ is sufficiently large, with Lemma 4.3 and Lemma 4.4 we have the estimate
$\#\mathcal{R}_{\theta,1}(dX^{1/2},X)$ $\geq\sum_{\begin{subarray}{c}2\leq\alpha\leq(c/4)X^{1/6}\\ \alpha\equiv 0\pmod{2}\end{subarray}}\left(\sum_{\begin{subarray}{c}1\leq\beta\leq\alpha\\ \gcd(\alpha,\beta)=1\end{subarray}}\#\mathbf{S}_{c}(\alpha,\beta,X)\right)$
$\geq\sum_{\begin{subarray}{c}2\leq\alpha\leq(c/4)X^{1/6}\\ \alpha\equiv 0\pmod{2}\end{subarray}}\left(\sum_{\begin{subarray}{c}1\leq\beta\leq\alpha\\ \gcd(\alpha,\beta)=1\end{subarray}}\frac{c\cdot X^{1/6}}{4\alpha}\right)$
$\gg\sum_{\begin{subarray}{c}2\leq\alpha\leq(c/4)X^{1/6}\\ \alpha\equiv 0\pmod{2}\end{subarray}}X^{1/6-\varepsilon}$
$\gg X^{1/3-\varepsilon}.$
Thus we prove that for sufficiently large $X$ we have
$\#\mathcal{R}_{\theta,1}(X,X)\geq\#\mathcal{R}_{\theta,1}(dX^{1/2},X)\gg X^{1/3-o(1)},$
which proves Lemma 1.5.
5. Further Remarks
Theorem 1.1 concerns itself with $25\%$ of degree $3$ monic irreducible polynomials. This limitation arises because we only transferred lower bounds from $\mathcal{R}_{\theta,1}(\cdot,\cdot)$ to $B_{\mathbf{p}}(\cdot)$. It is important to note that $\mathcal{R}_{\theta,1}(\cdot,\cdot)$ can be empty, for example, if $a_{1}a_{2}-a_{0}$ is even. To cover all cases of irreducible monic polynomials of degree $3$, we expect one would need to use sets $\mathcal{R}_{\theta,2^{t}}(\cdot,\cdot)$ for a suitable integer $t$ and employ enhanced parity arguments; however, we do not pursue these additional complications here. Furthermore, our arguments should generalize to study integer points on the Châtelet surface $y^{2}+nz^{2}=\mathbf{p}(x)$. Here our theory is stronger when $n\in\mathbb{N}$, as our unit arguments can be transferred quite readily, but when $n$ is negative, most of our theory breaks down (with the exception of an analogue for Lemma 1.3). This is because our arguments, even after suitable modification, do not provide sufficient information about the density of units in the ring $\mathcal{O}_{{\mathbb{Q}}(\theta,\sqrt{-n})}$.
6. Acknowledgements
The author would like to thank Igor Shparlinski for pointing out a result by Kronecker, which states that if a monic irreducible polynomial has all its roots on the unit circle then it is a cyclotomic polynomial. This research was supported by the Australian Government Research Training Program (RTP) Scholarship and a top-up scholarship from the University of New South Wales, both of which were instrumental in this work.